# A Web-based MPI Simulation Tool

### v1

@@ 1.2 Image Reconstruction
The MPI system is a linear-shift invariant system [5]. Hence, the $n$-dimensional signal vector $y$ stemming from an arbitrary distribution of particles $\rho$ can be understood as a weighted sum of signals from $\delta$-samples placed at different positions within the FOV that has been discretized into $m$ voxels. The weights are proportional to the number of particles at that position, and can be determined from a simple matrix equation:
$S\rho=y+\epsilon,~{}\mathrm{with}~{}S\in\mathbb{R}^{m\times n},\rho\in\mathbb{R}^{m},y\in\mathbb{R}^{n}.$ (6)
The matrix $S$ is called the system matrix, its columns correspond to the measured $n$-dimensional signal from a $\delta$-sample put at the $m$ different positions within the FOV, the term $\epsilon$ denotes possible noise in the measurement. In general $m<n$, which means we are faced with an overdetermined system, so the inverse of the matrix $S$ can not be defined in a straight-forward manner, and instead of solving for $\rho$ directly we need to consider the optimization problem [7]:
$\hat{\rho}=\underset{\rho\in\mathbb{R}^{m}}{\mathrm{argmin}}\|S\rho-(y+\epsilon)\|.$ (7)
@@
$\hat{\rho}=\underset{\rho\in\mathbb{R}^{m}}{\mathrm{argmin}}\left\{\|S\rho-(y+\epsilon)\|+\lambda_{r}\|\rho\|\right\}.$ (8)

### v2

@@ 1.2 Image Reconstruction
The MPI system is a linear-shift invariant system [5]. Hence, the $m$-dimensional signal vector $y$ stemming from an arbitrary distribution of particles $\rho$ can be understood as a weighted sum of signals from $\delta$-samples placed at different positions within the FOV that has been discretized into $n$ voxels. The weights are proportional to the number of particles at that position, and can be determined from a simple matrix equation:
$S\rho=y+\epsilon,~{}\mathrm{with}~{}S\in\mathbb{R}^{m\times n},\rho\in\mathbb{R}^{n},y\in\mathbb{R}^{m}.$ (6)
The matrix $S$ is called the system matrix, its columns correspond to the measured $m$-dimensional signal from a $\delta$-sample put at the $n$ different positions within the FOV, the term $\epsilon$ denotes possible noise in the measurement. In general $n<m$, which means we are faced with an overdetermined system, so the inverse of the matrix $S$ can not be defined in a straight-forward manner, and instead of solving for $\rho$ directly we need to consider the optimization problem [7]:
$\hat{\rho}=\underset{\rho\in\mathbb{R}^{n}}{\mathrm{argmin}}\|S\rho-(y+\epsilon)\|.$ (7)
@@
$\hat{\rho}=\underset{\rho\in\mathbb{R}^{n}}{\mathrm{argmin}}\left\{\|S\rho-(y+\epsilon)\|+\lambda_{r}\|\rho\|\right\}.$ (8)